# ELO-Rated Sequence Rewards: Advancing Reinforcement Learning Models

### v1

@@ 4 Experiments
In this section, we conduct experiments on a suite of Atari benchmark tasks to investigate the empirical performance of our proposed method. We compare our method with several baseline algorithms in the literature, and conduct an ablation study.
4.1 Performance in Atari Benchmark with episodic preference
4.1.1 Setups
To evaluate the behavior of ERRL, we compared its performance with several based algorithms across four Atari tasks. In ordinal based RL, all trajectories record some key events for expert judgment instead of returning a numerical reward. These tasks are long-term feedback problems (up to 5000 steps). We use the open source RL framework RLLib [24] as the baseline, which is an efficient parallel RL framework and adopts the standard preprocessing of Deepmind [7]. All results have been fine-tuned, and detailed settings can be found in Appendix B.
Since our experiment is based on ordinal RL, how to generate the ordinal model is crucial to the learning results. Here we take Pong as an example to illustrate the ordinal production process.
Pong is a game that simulates table tennis (ping-pong). In normal mode, if you shot the ball out of the opponent’s baseline, you will get 1 point; if you don’t catch the opponent’s shot, you will deduct 1 point. The player who reaches 21 points first wins the game. The final trajectory reward is $\mathcal{R}(\sigma)$, its own score minus the opponent’s score. And our evaluation criteria for trajectory $\sigma^{A}$ and trajectory $\sigma^{B}$ are:
@@ 4.1.1 Setups
$k$ indicates the length of the trajectory. In order to avoid the trajectory being too long, the game automatically terminates after the step reaches 5000.
The rationale behind this design is twofold: if defeating your opponent directly is not feasible, the strategy shifts towards dragging your opponent into extended play, aiming to hit as many of the opponent’s balls as possible. Conversely, when you can defeat your opponent, the approach is to leverage any means necessary to secure the win. So there are no more restrictions. For more game settings please refer to Appendix B.
4.1.2 Performance in Atari Benchmark
@@
A comparison of overall performance, as shown in Figure 2, shows that ERRL generally outperforms the baseline algorithms on these several problems. Our method can achieve higher sample efficiency and generate better policies after convergence. Because the returns are sparse, it is difficult for the original PPO to learn useful information. As for the previous reward decomposition function, it is mainly aimed at the decomposition of numerical reward. The reward obtained directly from the conversion of ordinal reward will increase the fluctuation of training. When the agent employs PPO for training with the reward conversion based on ELO Rating, it can learn an effective policy. This is because the ELO rating tends to amplify the performance gap between trajectories. In the long-term RL problems, ELO ratings are more conducive to highlighting the pros and cons of trajectories than the compact real rewards. So that the agent can find useful information.
@@ 4.2 Ablation Study
We study the value of $\eta$ in the ERRL algorithm under the condition of a fixed learning rate of $1\times 10^{-4}$. For the value of $K$, we refer to the original design of ELO rating to fix $K=0.04\times\eta$.
As described in Section 3, $\eta$ controls the fluctuation range of the ELO rating, smaller $\eta$ represents greater randomness of the zero-sum game, larger training fluctuations, but training faster; larger $\eta$ represents less randomness, and reduced training fluctuations, but training slower. So we need a suitable $\eta$ to balance training speed and fluctuation.
@@
We perform an ablation study on the hyperparameter $\eta$ that represents the range of fluctuations in the ELO rating. Considering that the time steps of the game will change with the progress of the game, we decided to relating $\eta$ with the time-steps of the trajectory, that is, the $\eta$ actually used in the ELO Rating calculation is $\eta_{\text{elo}}=\eta_{\text{our}}\times\overline{k_{X}}$. For the convenience of writing, we will use $\eta$ to represent $\eta_{\text{our}}$ in the following text if there is no special emphasis.
We evaluate ERRL with a set of choices of hyperparameter $\eta\in\{0.1,0.05,0.01,0.005,0.001\}$. The experimental results shown in Figure 3 show that this hyperparameter has a great impact on RL training. A larger $\eta$ will cause the training fluctuation to be too large and directly collapse, while a smaller $\eta$ will cause the training to be too slow and easily lead to over fitting. An acceptable range of $\eta$ is in interval $[0.05,0.005]$.
From the time-steps of the entire trajectory, regardless of the size of $\eta$, the final time-steps converge to 2100$\sim$2400. It can be seen that under the same preference, if the training can converge, the choice of $\eta$ does not change the final strategies, and we will discuss the relationship between strategy selection and expert preference in Appendix C.
@@ 5 Related Work
RL without Environmental Reward. While RL typically relies on environmental reward, several studies have explored RL without direct environmental rewards. This research can be broadly categorized into two approaches. The first approach focuses on capturing essential information within trajectories and using these differences to define learning rewards Eysenbach et al. [28] proposed a pre-training method to train the discriminator to identify which agent generated the trajectory. the agent always hopes that the discriminator can easily identify itself, so the agent always tends to find a unique state. Hu et al. [26] uses multidimensional features to determine the Pareto frontier. Only the agents at the Pareto frontier will perform the next training. This approach enables game designers to create a variety of agents with distinct styles, but this type of method is generally only suitable for increasing the diversity of agents. Another idea is to use human embeddings in RL, such as PBRL [16] introduced in Section 2.4. In addition, Kimin et al. [13] proposed the method of PEBBLE, which has demonstrated exceptional outcomes. .
Sparse Rewards RL. Sparse rewards significantly hinder the application of RL [11, 8], with numerous studies addressing this challenge. A straightforward solution is the integration of human experience [1, 4]. As long as the human benefit is designed perfectly enough, this approach can lead to successful outcomes. However, relying heavily on human expertise can be expensive and may inadvertently limit the capabilities of the RL agent [1, 4, 5]. Another approach is reward decomposition, which decomposes the total reward into individual state-action pairs [15, 20]. usually this method requires reward shaping to achieve good results. Marcin et al. [10] also achieves good results by reusing historical information, but this only works for tasks with simple and clear states.
@@ 6 Conclusion and Future Work
We propose a new preference-based RL algorithm ERRL, which can achieve good performance in the 5000-step long-tern trajectory feedback RL problem. ERRL abandons the classical numerical reward, and estimates the ordinal reward of each trajectory through the expert’s preference and the ELO rating method. In addition, a new reward decomposition method is adopted to reduce the fluctuation of training in ERRL setting. Experiments show that ERRL outperforms previous methods in solution quality. In future work, we will explore the theoretical basis for combining ELO rating with RL, conduct more detailed experiments, discuss the effectiveness of our algorithm when it is difficult for experts to make judgments, and explore the feasibility of ELO rating in the remaining areas. Such as reward shaping  [27], games with hard-to-determine payoffs (ie. Texas hold’em and Hanabi) [25], or research on the diversity of RL agents [26, 28].
@@
According to maximum likelihood estimation, the above problem can be transformed into an optimization problem. The current formula  17 takes the maximum value of $x_{t}$ as the most likely update range:
@@
Taking the logarithm of the original formula 17, there are:
@@
Our open source main experimental code and details can be found in https://github.com/Zealoter/ERRL, as our experiments are still ongoing, we will gradually improve our code and experimental details.

### v2

@@ 4 Experiments
In this section, we conduct experiments on a suite of Atari benchmark tasks to evaluate the empirical performance of our proposed method. We compare ERRL with several state-of-the-art baselines and perform ablation studies to analyze its components.
4.1 Experimental Setups
To assess ERRL’s behavior, we compare its performance against multiple baseline algorithms across four Atari tasks, all featuring long-term feedback (up to 5000 steps). In these ordinal-based RL settings, trajectories record key events for expert preference judgments rather than returning numerical rewards. We use the open-source RL framework RLLib [24] as the baseline, which adopts DeepMind’s standard preprocessing pipeline [7] and enables efficient parallel training. All results are fine-tuned, with detailed configurations provided in Appendix Appendix B.
As our experiments rely on ordinal RL, generating the ordinal preference model is critical for learning outcomes. We illustrate the ordinal generation process using Pong as a case study.
[…]
@@ 4.1 Experimental Setups
    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$, then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$; otherwise, proceed to the next criterion.
@@
    If both $\mathcal{R}(\sigma^{A})>0$ and $\mathcal{R}(\sigma^{B})>0$, then $\sigma^{A}\approx\sigma^{B}$; otherwise, proceed to the third criterion.
@@
    If the trajectory length $k_{A}>k_{B}$ or $k_{A}<k_{B}$, then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$; otherwise, $\sigma^{A}\approx\sigma^{B}$.
Here, $k$ denotes the trajectory length, and the game terminates automatically after 5000 steps to prevent excessively long trajectories.
This design follows two principles: when direct victory is unfeasible, the strategy shifts to prolonged play, aiming to maximize ball returns; when victory is achievable, any means to secure the win are prioritized. No additional constraints are imposed. For detailed game configurations, refer to Appendix Appendix B.
4.2 Performance in Atari Benchmark
@@
The overall performance comparison in Figure 2 demonstrates that ERRL consistently outperforms baseline algorithms across these tasks, achieving higher sample efficiency and converging to superior policies. Due to the sparse nature of rewards, vanilla PPO struggles to extract meaningful information. Existing reward decomposition methods, designed for numerical rewards, introduce instability when applied to ordinal preferences. In contrast, ERRL leverages ELO ratings to convert ordinal feedback into scalable rewards, amplifying performance differences between trajectories. This approach proves particularly effective in long-horizon RL problems, where ELO ratings more effectively highlight trajectory quality than raw numerical rewards, enabling the agent to identify optimal strategies.
@@ 4.3 Ablation Study
We study the impact of $\eta$ in the ERRL algorithm under a fixed learning rate of $1\times 10^{-4}$. For the parameter $K$, we follow the original ELO design by setting $K=0.04\times\eta$.
As discussed in Section 3, $\eta$ regulates the fluctuation range of ELO ratings: smaller $\eta$ introduces higher zero-sum game randomness, leading to larger training volatility but faster convergence; larger $\eta$ reduces randomness, decreasing training fluctuations but slowing down learning. Thus, an optimal $\eta$ is needed to balance training speed and stability.
@@
We conduct an ablation study on the hyperparameter $\eta$, which controls the fluctuation range of the ELO rating. Given that game time steps vary during gameplay, we link $\eta$ to the trajectory length, such that the effective $\eta$ value used in ELO rating calculations is $\eta_{\text{elo}}=\eta_{\text{our}}\times\overline{k_{X}}$. For simplicity, we use $\eta$ to denote $\eta_{\text{our}}$ throughout the remainder of the text unless otherwise specified.
We evaluate ERRL with $\eta\in\{0.1,0.05,0.01,0.005,0.001\}$. As illustrated in Figure 3, this hyperparameter significantly impacts RL training. Larger $\eta$ values lead to excessive training volatility and potential model collapse, whereas smaller $\eta$ values slow down convergence and increase the risk of overfitting. An optimal range for $\eta$ is identified as $[0.05,0.005]$.
Across all trajectories, regardless of $\eta$’s magnitude, the final time steps converge within the range of 2100–2400. This indicates that, under consistent expert preferences, the choice of $\eta$ does not alter the learned optimal strategies. We further discuss the relationship between strategy selection and expert preferences in Appendix Appendix C.
@@ 5 Related Work
RL without Environmental Reward.
While traditional RL relies on environmental rewards, recent studies have explored reward-free RL frameworks, which can be categorized into two main approaches. The first approach focuses on extracting discriminative information from trajectories to define surrogate rewards. Eysenbach et al. [28] proposed a pre-training method where a discriminator is trained to identify the agent that generated a trajectory. The agent aims to produce trajectories that are easily distinguishable, driving it to seek unique states. Hu et al. [26] uses multidimensional features to determine the Pareto frontier, with only frontier agents proceeding to the next training phase. This approach enhances agent diversity but is typically limited to diversity-driven tasks. The second approach integrates human preferences into RL, such as Preference-Based RL (PBRL) [16] discussed in Section 2.4. Additionally, Kimin et al. [13] proposed PEBBLE, which has shown remarkable performance in reward-free settings.
Sparse Rewards RL.
Sparse rewards remain a critical challenge for RL applications [11, 8], with research addressing this through multiple strategies. A direct solution is to incorporate human expertise [1, 4], which can yield successful results if expert knowledge is well-codified. However, heavy reliance on human input is costly and may inadvertently constrain agent capabilities [1, 4, 5]. Another approach is reward decomposition, which distributes cumulative rewards to individual state-action pairs [15, 20]. This often requires reward shaping to achieve effectiveness. Marcin et al. [10] leverages historical information reuse for sparse reward tasks, though this is limited to domains with simple, well-defined states.
@@ 6 Conclusion and Future Work
We propose a novel preference-based RL algorithm, ERRL (ELO-Rating based Reinforcement Learning), which demonstrates superior performance in 5000-step long-term trajectory-feedback RL tasks. ERRL eschews classical numerical rewards, instead estimating ordinal rewards for each trajectory via expert preferences and the ELO rating system. Additionally, a new reward decomposition method is introduced to mitigate training volatility under the ERRL framework. Experimental results show that ERRL outperforms state-of-the-art methods in solution quality. For future work, we will explore the theoretical foundations of integrating ELO rating with RL, conduct more detailed ablation studies, investigate algorithm effectiveness under ambiguous expert preferences, and explore the feasibility of ELO rating in extended domains including reward shaping [27], games with complex payoffs (e.g., Texas hold’em and Hanabi) [25], and research on RL agent diversity [26, 28].
@@
According to maximum likelihood estimation, the above problem can be transformed into an optimization problem. The current formula  18 takes the maximum value of $x_{t}$ as the most likely update range:
@@
Taking the logarithm of the original formula 18, there are:
@@
Our open source main experimental code and details can be found in Github, as our experiments are still ongoing, we will gradually improve our code and experimental details.